OpenStudy (anonymous):

Solve the system of equations. x + 3y = -1 2x + 2y = 6 (-4, 1) (2, -1) (4, -1) (5, -2)

OpenStudy (anonymous):

y=-2

OpenStudy (anonymous):

really need help

OpenStudy (anonymous):

so answer is D

OpenStudy (anonymous):

(5, -2)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

please help me on one more?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Solve the following system of equations: -2x + y = 1 -4x + y = -1 (3, 1) (-1, 3) (-1, -3) (1, 3)

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

so D

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you're welcome